OpenStudy (anonymous):

how dou you factor 3t^2+5t-1

OpenStudy (anonymous):

This trinomial can not be factored. Why? Because the Discriminant is not a perfect square. D = b^2 - 4ac = 25 + 12 = 37

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

3t^2+5t-1 use quadratic formula. \[(-b+/-\sqrt{b ^{2}-4ac})/2a\]

OpenStudy (anonymous):

so it would be -5+-square root of 25+12/6 =(-5+-sqrt of 37)/6 and solve